# Segment Anything Meets Point Tracking

### v1

@@ Appendix A Point Tracking Reinitialization
In our SAM-PT-reinit method, we introduce a reinitialization strategy. Here, the point tracker begins anew after every $h$ frames, where $h$ represents a pre-set tracking horizon (e.g., 8 frames), or is dynamically determined based on SAM’s mask predictions for each timestep within the horizon (e.g., using most-similar-mask-area heuristics). Upon reaching this horizon, the query points given to the tracker are reinitialized according to the mask prediction SAM outputted at the horizon frame. While this method may increase the computational load (especially if some of SAM’s computed masks are disregarded), it demonstrates substantial performance improvement in demanding video sequences, such as those in the MOSE dataset.
@@
We explored four reinitialization methods, each varying in how they compute the value of $h$:
@@
    Reinit-on-Horizon-and-Sync-Masks: This straightforward variant reinitializes points after a fixed number of frames (e.g., every 8 frames). However, it may stumble if the mask is absent at the reinitialization timestep. Despite this potential pitfall, it operates at the same speed as methods that do not employ reinitialization.
@@
    Reinit-at-Median-of-Area-Diff: In this variant, the tracker outputs trajectory points for each frame within the horizon, and SAM predicts masks based on these trajectories. Reinitialization happens at the frame within the horizon that has the mean mask area among the non-empty masks predicted by SAM. Notably, this approach may be significantly slower than methods without reinitialization, as it may reject several SAM masks (e.g., out of 8 computed masks, reinitialization might occur on the second one, necessitating recomputation of the remaining 6 masks in the next step).
@@
    Reinit-on-Similar-Mask-Area: This method triggers reinitialization when the mask area is similar to the initial mask area, causing it to be several times slower than methods without reinitialization.
@@
    Reinit-on-Similar-Mask-Area-and-Sync-Masks: This variant reinitializes when the mask area for all masks in the batch is similar to the initial mask areas, synchronizing the masks to be tracked from the same timestep. This synchronization allows for the use of negative points from other masks when querying SAM, but it also runs several times slower than methods without reinitialization.
@@
From our investigations, we found the (A) Reinit-on-Horizon-and-Sync-Masks strategy to be the most effective, as indicated by its superior performance on the DAVIS $2017$ validation subset. The choice of reinitialization method may depend on the specific validation subset and the degree of hyperparameter tuning involved. Note that we have always used reinitialization along with negative points.
A.1 Computational Cost and Speed Optimization
The introduction of reinitialization in SAM-PT-reinit comes with a trade-off: it slows down the inference speed by a factor of 2 to 8, depending on the reinitialization method and parameters used. The major bottleneck is the invocation of SAM’s backbone for each video frame. We propose caching the backbone outputs for unprocessed video frames as a possible solution to mitigate this slowdown. This strategy requires storing embeddings for all video frames in the working memory but offers the potential for significant speedup, particularly useful for applications requiring faster inference.
Appendix B More DAVIS Subsets
We report results on DAVIS $2016$ Validation and DAVIS $2017$ Test-dev in Tab. 7.
Table 7: Quantitative results in semi-supervised video object segmentation on additional subsets of DAVIS. ♠: our reproduced result using the official code of [45].
DAVIS $\mathbf{2016}$ Validation [27] DAVIS $\mathbf{2017}$ Test-dev [27]
[…]
PerSAM-F ♠ [45] $74.8$ $74.5$ $75.0$ $47.6$ $45.5$ $49.7$
SegGPT [36] $82.3$ $81.8$ $82.8$- - -
SAM-PT $\mathbf{83.1}\pm 1.5$ $\mathbf{83.0}\pm 0.8$ $\mathbf{83.0}\pm 1.1$ $\mathbf{62.7}\pm 0.5$ $\mathbf{59.4}\pm 0.6$ $\mathbf{66.1}\pm 0.4$
SAM-PT-reinit $80.2\pm 0.6$ $80.3\pm 0.6$ $80.0\pm 0.6$ $61.5\pm 1.1$ $59.3\pm 1.0$ $63.8\pm 1.2$
Appendix C Per-sequence DAVIS 2017 Validation Results
See figure Fig. 10 for per-sequence DAVIS 2017 Validation results. For exact numbers and tables, check our GitHub experiment summaries or the Wandb project.
(a) number of positive points per mask
(b) point selection methods
(c) point tracking module
(d) number of negative points per mask
(e) number of iterative refinement iterations
(f) patch similarity filtering threshold
Figure 10: Per-sequence Ablation Results. The charts plot the median J-Mean result of different ablation result experiments along with $95\%$ bootstrapped confidence intervals, based on 1000 bootstrap samples.

### v2

@@ Appendix G Point Tracking Reinitialization
In our method, we introduce an optional reinitialization strategy. Here, the point tracker begins anew every $h$ frames, where $h$ represents a pre-set tracking horizon (e.g., $8$ frames), or is dynamically determined based on SAM’s mask predictions for each timestep within the horizon (e.g., using most-similar-mask-area heuristics). Upon reaching this horizon, the query points given to the tracker are reinitialized according to the mask prediction SAM outputted at the horizon frame. While this method may increase the computational load, it shows performance improvement when using the PIPS [16] point tracker in demanding video sequences, such as those in the MOSE dataset. However, our studies also suggested that the proposed reinitialization strategies hurt performance when using CoTracker as the point tracker. We primarily designed the reinitialization variants to address some of the failure cases of PIPS such as the common case of points being wrongly predicted to be off the target object, being predicted on the background instead, but this does not work as well for other point trackers such as CoTracker that are more robust to such failures.
@@
We explored four reinitialization strategies, each varying in how they compute the value of $h$:
@@
    Reinit-on-Horizon-and-Sync-Masks: This straightforward variant reinitializes points after a fixed number of frames (e.g., every 8 frames). However, it may stumble if the mask is absent at the reinitialization timestep.
@@
    Reinit-at-Median-of-Area-Diff: In this variant, the tracker outputs trajectory points for each frame within the horizon, and SAM predicts masks based on these trajectories. Reinitialization happens at the frame within the horizon that has the mean mask area among the non-empty masks predicted by SAM.
@@
    Reinit-on-Similar-Mask-Area: This method triggers reinitialization when the mask area is similar to the initial mask area.
@@
    Reinit-on-Similar-Mask-Area-and-Sync-Masks: This variant reinitializes when the mask area for all masks in the batch is similar to the initial mask areas, synchronizing the masks to be tracked from the same timestep. This synchronization allows for the use of negative points from other masks when querying SAM.
@@
From our ablation investigations, we found the (A) Reinit-on-Horizon-and-Sync-Masks strategy to be effective with PIPS as the point tracker and (B) Reinit-at-Median-of-Area-Diff with CoTracker. Besides the point tracker used, the choice of reinitialization method may depend on the specific validation subset and the degree of hyperparameter tuning involved. Note that we always use reinitialization along with negative points.
Appendix H Additional Qualitative Results
We show additional visualizations on DAVIS videos in Fig. 8. We show failure cases on clips from the anime-influenced series “Avatar: The Last Airbender” in Fig. 9.
Appendix I VOS Evaluation Details
Appendix J VIS Evaluation Details
For evaluating our method on the VIS task, we leverage SAM’s automatic mask generation capacity to generate up to 100 mask proposals for the initial frame. We use the same initial masks to assess TAM [46] for a fair comparison. Our current approach does not generate new proposals in subsequent frames, thus it cannot detect new objects appearing after the first frame, unlike fully-fledged VIS methods. However, this setup allows for a straightforward comparison of zero-shot capabilities in mask propagation from the initial frame.
Appendix K BDD100K VOS Dataset Creation
BDD100K is a large open driving video dataset with 100K videos and 10 tasks to evaluate the progress of image recognition algorithms in autonomous driving. It includes a variety of geographic, environmental, and weather conditions. We convert its annotations from the Multi-Object Tracking and Segmentation (MOTS) section into semi-supervised VOS annotations. This section has $154$, $32$, and $37$ videos for train, validation, and test sets, totaling $25$K instances and $480$K masks.
To convert the annotations, we take the ground truth mask of the first appearance of each object in the videos. This mask will be given as input to semi-supervised VOS methods which then need to predict the masks for the remaining video frames. In converting the validation subset of $32$ videos and $4566$ object tracks, we create $61$ semi-supervised VOS datapoints of up to $100$ objects per video. We limit the number of objects per video for implementation simplicity since most VOS methods expect a small number of objects per video. For example, in the DAVIS validation subset, the maximum number of objects per video is 5. During the conversion, we additionally remove the instances marked as “ignored” or “crowd” in the MOTS annotations.
Appendix L Interactive Point-Based Video Segmentation Details
Interactive point-based video segmentation aims to refine segmentation masks with minimal user input while optimizing the Intersection over Union (IoU) with the ground truth. To evaluate the responsiveness of SAM-PT to simulated human input, we benchmark against a SAM-only baseline (Algorithm 1) and compare it with both an online (Algorithm 2) and offline (Algorithm 3) interactive SAM-PT method:
  1. 1.
    Non-tracking method (SAM only): As a baseline, this method mimics user interactions by selecting points on each frame without any point tracking, effectively emulating the process of manual, frame-by-frame annotation using the standalone SAM model.
  2. 2.
    Online method: This approach models a user going through the video sequentially a single time, making corrective per-frame interactions to achieve an IoU of at least 95% with the ground truth mask for each frame. These corrections include the addition or removal of points and are propagated to subsequent frames via point tracking. The user may opt to skip frames that already appear to be well-annotated or that cannot be annotated sufficiently well within the available interaction budget.
  3. 3.
    Offline method (Checkpoint Strategy): This method takes a multi-pass approach, incrementally aiming for higher IoU thresholds with each pass through the video. Checkpoints are saved after each pass, and the latest checkpoint given the interaction budget is returned, allowing for a more global optimization approach.
Interactions are defined as the act of adding or removing a point and are executed as described in Algorithm 4. To maintain simplicity in our evaluation, skipping frames while progressing through a video is not counted as an interaction.
Algorithm 1 Non-tracking (SAM only) Method.
1:   Input: rgbs, gt_masks, int_per_frame
2:   Output: pred_masks
3:   point_memory $\leftarrow$ emptyPointMemory()
4:   for i in frames do
5:      pmf $\leftarrow$ emptyFramePointMemory()
6:      for j in int_per_frame do
7:         m $\leftarrow$ predMask(rgbs[i], pmf)
8:         performInt(i, m, gt_masks[i], pmf)
9:      end for
10:      addToPointMemory(point_memory, i, pmf)
11:   end for
12:   return  predAllMasks(rgbs, point_memory)
Algorithm 2 Online Method.
1:   Input: rgbs, gt_masks
2:   Output: pred_masks
3:   max_int $\leftarrow 300$
4:   max_int_per_frame $\leftarrow 3$
5:   threshold $\leftarrow 0.95$
6:   point_memory $\leftarrow$ selectFirstPoint(gt_masks[0])
7:   max_int $\leftarrow$ max_int - 1
8:   for i in frames do
9:      m $\leftarrow$ predMask(rgbs[i], point_memory)
10:      IoU $\leftarrow$ calculateIoU(m, gt_masks[i])
11:      if IoU $\geq$ threshold then
12:         continue
13:      end if
14:      performInt(i, m, gt_masks[i], point_memory)
15:      max_int $\leftarrow$ max_int - 1
16:      if max_int $\leq 0$ then
17:         break
18:      end if
19:   end for
20:   return  predAllMasks(rgbs, point_memory)
Algorithm 3 Offline Method (Checkpoint Strategy).
1:   Input: rgbs, gt_masks
2:   Output: pred_masks
3:   max_int $\leftarrow 300$
4:   max_int_per_frame $\leftarrow 3$
5:   int_iou_thresholds $\leftarrow[0.10,0.20,\ldots,0.95]$
6:   point_memory $\leftarrow$ selectFirstPoint(gt_masks[0])
7:   max_int $\leftarrow$ max_int - 1
8:   best_ckpt $\leftarrow$ predAllMasks(rgbs, point_memory)
9:   for threshold in int_iou_thresholds do
10:      for i in frames do
11:         for j in max_int_per_frame do
12:            m $\leftarrow$ predMask(rgbs[i], point_memory)
13:            IoU $\leftarrow$ calculateIoU(m, gt_masks[i])
14:            if IoU $\geq$ threshold then
15:               break
16:            end if
17:            performInt(i, m, gt_masks[i], point_memory)
18:            max_int $\leftarrow$ max_int - 1
19:         end for
20:      end for
21:      if max_int $\geq 0$ then
22:         best_ckpt $\leftarrow$ predAllMasks(rgbs, point_memory)
23:      else
24:         break
25:      end if
26:   end for
27:   return  best_ckpt
Algorithm 4 Perform Interaction.
1:   Input: frame_idx, m, gt_m, point_memory
2:   i $\leftarrow$ frame_idx
3:   tp_mask $\leftarrow$ m $\land$ gt_m
4:   tn_mask $\leftarrow$ $\lnot$¬m $\land$ $\lnot$¬gt_m
5:   fp_mask $\leftarrow$ m $\land$ $\lnot$¬gt_m
6:   fn_mask $\leftarrow$ $\lnot$¬m $\land$ gt_m
7:   pos_points $\leftarrow$ getPosPoints(point_memory, i)
8:   neg_points $\leftarrow$ getNegPoints(point_memory, i)
9:   if any neg_points in fn_mask then
10:      p $\leftarrow$ firstIncorrectNegPoint(fn_mask, neg_points)
11:      removePointAndItsFuture(point_memory, i, p)
12:   else if any pos_points in fp_mask then
13:      p $\leftarrow$ firstIncorrectPosPoint(fp_mask, pos_points)
14:      removePointAndItsFuture(point_memory, i, p)
15:   else
16:      if sum(fn_mask) $>$ sum(fp_mask) then
17:         [x, y] $\leftarrow$ extractPoint(fn_mask)
18:         lbl $\leftarrow$ positive
19:      else
20:         [x, y] $\leftarrow$ extractPoint(fp_mask)
21:         lbl $\leftarrow$ negative
22:      end if
23:      addToFrameAndFuture(point_memory, i, lbl, x, y)
24:   end if
(a) Successful cases for SAM-PT.
(b) Failure cases for SAM-PT.
Figure 8: Additional visualization of SAM-PT on videos from the DAVIS 2017 [30] validation subset, including (a) successful cases and (b) failure cases. Circles denote positive points and crosses denote negative points. Red symbols indicate occlusion prediction.